OpenStudy (shanan1234):

A computer manufacturer built a new facility for assembling computers. There were construction and new equipment costs. The company paid for these costs and made combined profits of $80 million after 4 years, while profits increased $60 million per year. Select the correct graph of this function.

OpenStudy (wanell):

Given a point (4,80) and a slope, m = 60, these can be used to formulate the equation of the graph. Using the point-slope formula: y - y1 = m(x - x1) y - 80 = 60(x - 4) y - 80 = 60x - 240 y = 60x -160 Taking the x and y-intercepts: At y = 0 0 = 60x - 160 x = 2.7 At x = 0 y = 60(0) - 160 y = -160 Therefore, you could state that there is a graph of a line that has a y intercept that is approximately negative $160 million and an x intercept that is approximately 2.7 years.

OpenStudy (shanan1234):

My choice options are limited and i don't exactly know how to get the graphs so that I can show you what they are. Could I explain what they look like to you?

OpenStudy (shanan1234):

I figured it out thanks so much for your help! :)